# CarbonCall: Sustainability-Aware Function Calling for Large Language Models on Edge Devices

### v1

@@ I Introduction
Edge systems have limited power and computing resources, making LLM execution inefficient [3]. Function calling increases energy demands further, as each query triggers additional inference cycles, leading to higher latency and power consumption. Unlike cloud servers, which optimize workloads using large-scale energy-efficient computing, edge devices rely on localized, often carbon-intensive power sources that vary by region [4, 5]. As agentic LLMs become more common in real-time applications [2], their energy footprint raises sustainability concerns, requiring efficient execution without compromising responsiveness. Existing optimizations, such as replacing large models (e.g., 70B, 400B parameters) with smaller LLMs (e.g., 1.5B, 3.8B, 7B, 8B), improve efficiency but focus primarily on performance, often neglecting sustainability [3, 1]. Addressing these challenges requires sustainability-aware function-calling strategies that reduce power consumption while maintaining responsiveness.
@@ II Related Work
Existing work tackles two important focus areas of this paper: (i) Edge LLM inference and function calling methods; (i) Carbon aware execution strategies. To enhance execution efficiency, EdgeLLM [9] proposes a confidence mechanism to prune unlikely branches in the token tree and eliminate wasteful computations. Although speculative execution improves efficiency in general, its impact on function calling remains underexplored. STE [10] introduces a trial-and-error fine-tuning approach for faster LLM inference. However, the high cost of training domain-specific LLMs for different tasks and device characteristics makes this approach impractical for edge deployment. ToolLLM [11] employs a tree-based selection mechanism to reduce the number of tools required for function execution. Unfortunately, in cases when the entire toolset needs to be enumerated, ToolLLM suffers from increased latency and energy consumption. Moreover, Less-is-More [1] proposes an LLM-based tool recommender to identify and prefetch potential tools. However, this method relies on a three-level offline clustering mechanism, creating scalability issues. Recent work on efficient function-calling, such as Gorilla [2] and TinyAgent [12], leverages Retrieval-Augmented Generation (RAG) and a transformer-based classifier, respectively, to fetch relevant APIs. However, these techniques might lack generalizability across diverse function spaces, limiting their applicability to dynamic and sustainability-aware function calling at the edge.
@@ III-D Quantized LLMs for low-power modes
However, as quantization levels decrease, these models tend to exhibit lower function-calling accuracy [1]. Highly compressed models struggle with more complex tool selection tasks, leading to higher failure rates when identifying the correct API function. This highlights the importance of CarbonCall’s tool selection mechanism (Section III-B), which compensates for the loss in LLM accuracy by ensuring that only the most relevant tools are retrieved and passed to the model.
@@ IV Experimental results
We evaluated CarbonCall using the BFCL [19] and GeoEngine [20] benchmarks. Queries in BFCL require single function calls per sub-question, while GeoEngine involves multiple sequential function calls, making them challenging. To ensure a diverse difficulty distribution and realistic user-based input, we combined queries from both benchmarks. All experiments were conducted on an NVIDIA Jetson AGX Orin board.
@@
We compared CarbonCall against: (i) Defaultmethod, which invokes the LLMs with all the available functions and no other optimization; (i) Gorilla[2], which filters functions by directly matching the queries and the function descriptions; (i) Less-is-More (LiS) [1], which uses the LLM’s reasoning capabilities to generate its own set of functions per query; (i) LiS*, a modified version of LiS, which changes the operation modes according to the carbon intensity while running the same quantized variant. We evaluated our method by running experiments on all these approaches and comparing the normalized values of: (i) average latency (TNorm); (i) average power consumption (PNorm); (i) average token per second (TPSNorm); and (i) average carbon emission per query (CFNorm).
@@ VI Acknowledgement
This work is supported by grant NSF CCF 2324854.

### v2

@@ I Introduction
Edge systems have limited power and computing resources, making LLM execution inefficient [3, 4]. Function calling increases energy demands further, as each query triggers additional inference cycles, leading to higher latency and power consumption. Unlike cloud servers, which optimize workloads using large-scale energy-efficient computing, edge devices rely on localized, often carbon-intensive power sources that vary by region [5, 6]. As agentic LLMs become more common in real-time applications [2], their energy footprint raises sustainability concerns, requiring efficient execution without compromising responsiveness. Existing optimizations, such as replacing large models (e.g., 70B, 400B parameters) with smaller LLMs (e.g., 1.5B, 3.8B, 7B, 8B), improve efficiency but focus primarily on performance, often neglecting sustainability [3, 1]. Addressing these challenges requires sustainability-aware function-calling strategies that reduce power consumption while maintaining responsiveness.
@@ II Related Work
Existing work tackles two important focus areas of this paper: (i) Edge LLM inference and function calling methods; (i) Carbon aware execution strategies. To enhance execution efficiency, EdgeLLM [10] proposes a confidence mechanism to prune unlikely branches in the token tree and eliminate wasteful computations. Although speculative execution improves efficiency in general, its impact on function calling remains underexplored. To optimize function execution, caching-based approaches have been explored, such as query- and tool-caching [11], which store frequently used LLM responses to reduce redundant computation. STE [12] introduces a trial-and-error fine-tuning approach for faster LLM inference. However, the high cost of training domain-specific LLMs for different tasks and device characteristics makes this approach impractical for edge deployment. ToolLLM [13] employs a tree-based selection mechanism to reduce the number of tools required for function execution. Unfortunately, in cases when the entire toolset needs to be enumerated, ToolLLM suffers from increased latency and energy consumption. Moreover, Less-is-More [1] proposes an LLM-based tool recommender to identify and prefetch potential tools. However, this method relies on a three-level offline clustering mechanism, creating scalability issues. Recent work on efficient function-calling, such as Gorilla [2] and TinyAgent [14], leverages Retrieval-Augmented Generation (RAG) and a transformer-based classifier, respectively, to fetch relevant APIs. However, these techniques might lack generalizability across diverse function spaces, limiting their applicability to dynamic and sustainability-aware function calling at the edge.
@@ III-D Quantized LLMs for low-power modes
However, as quantization levels decrease, these models tend to exhibit lower function-calling accuracy [1]. Highly compressed models struggle with more complex tool selection tasks, leading to higher failure rates when identifying the correct API function. This highlights the importance of CarbonCall’s tool selection mechanism (Section III-B), which compensates for the loss in LLM accuracy by ensuring that only the most relevant tools are retrieved and passed to the model. By integrating quantized LLMs with robust tool selection, CarbonCall ensures that function calling remains effective even at lower power configurations.
@@ III-E Runtime
In summary, CarbonCall integrates real-time CI forecasts, adaptive power management, and quantized LLM transitions to optimize function execution for both sustainability and performance. By dynamically adjusting power thresholds and model variants, it effectively reduces the carbon footprint while maintaining a high-quality user experience in edge-based AI deployments, as shown in Section IV.
@@ IV Experimental results
We evaluated CarbonCall using the BFCL [21] and GeoEngine [11] benchmarks. Queries in BFCL require single function calls per sub-question, while GeoEngine involves multiple sequential function calls, making them challenging. To ensure a diverse difficulty distribution and realistic user-based input, we combined queries from both benchmarks. All experiments were conducted on an NVIDIA Jetson AGX Orin board. which offers (i) 2048-core NVIDIA Ampere architecture GPU with 64 Tensor Cores; (i) 12-core ARM Cortex-A78 CPU with 3 clusters of 4 cores each; and (i) 64GB LPDDR5 memory. With these specs, the Jetson AGX Orin can perform up to 270 TOPS, capped at 60W, making it a suitable platform to evaluate small-sized LLMs in edge settings.
@@
We compared CarbonCall against: (i) Defaultmethod, which invokes the LLMs with all the available functions and no other optimization; (i) Gorilla[2], which filters functions by directly matching the queries and the function descriptions; (i) Less-is-More (LiS) [1], which uses the LLM’s reasoning capabilities to generate its own set of functions per query; (i) LiS*, a modified version of LiS, which changes the operation modes according to the carbon intensity while running the same quantized variant.
We evaluated our method by running experiments on all these approaches and comparing the normalized values of: (i) average latency (TNorm); (i) average power consumption (PNorm); (i) average token per second (TPSNorm); and (i) average carbon emission per query (CFNorm).
@@ VI Acknowledgement
This work is supported by grant NSF CCF 2324854. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.